A ball weighing 500n is drop from a height of 30 meters find the kinetic energy of the ball just before hitting the ground.
1
Expert's answer
2020-10-09T07:03:17-0400

The ball hits the ground in the time given by the kinematic law:


"t = \sqrt{\dfrac{2h}{g}}"

where "h = 30m" is the height and "g = 9.81m\/s^2" is the gravitational acceleration.

The speed of the ball at this time:


"v = gt =g\sqrt{\dfrac{2h}{g}}= \sqrt{2gh}"

The mass of the ball (according to the second Newton's law):


"m = \dfrac{F}{g}"

where "F = 500N" is the weight of the ball.

By definition, the  kinetic energy is:


"K = \dfrac{mv^2}{2} = \dfrac{F}{g}\cdot \dfrac{\sqrt{2gh}^2}{2} = gFh^2\\
K = 9.81\times 500\times 30^2 = 4.4145 \times 10^6L = 4.4145\space MJ"

Answer. "4.4145\space MJ."